Solve the equation: 2x - 5 = -13

Answer by checkley75(3666)

2X-5=-13
2X=-13+5
2X=-8
X=-8/2
X=-4 ANSWER.
PROOF
2*-4-5=-13
-8-5=-13
-13=-13